# JUREX-4E: Juridical Expert-Annotated Four-Element Knowledge Base for Legal Reasoning

### v1

@@ 7.4 Results
The LCR results are shown in Table 4, where we can observe that:
FET Works Well in LCR.
The baseline model BGE achieves strong performance across most metrics compared to previous methods. Introducing the Four-Element Theory (FET) further improves its results, with relative MRR improvements of 11.11% for FETLLM and 11.89% for FETExpert_guided, indicating that introducing legal theory is effective.
Expert Knowledge is Necessary.
By leveraging external knowledge, FETExpert_guided achieves significant improvements across all of the metrics. Specifically, using expert-guided case four-elements (FETExpert_guided-base) outperforms LLM-generated case four-elements (FETLLM-base) by an average of 11.77% in MRR, demonstrating the critical role of expert knowledge in enhancing retrieval precision. A case study in Appendix G shows that the expert four-element for charges provide practical judgment points and key narratives (e.g., the special subject of the Crime of Embezzlement) that help the LLM focus on essential facts to analyze the case.
We also evaluated the FET method on the full set, as shown in Table 9 , and the results remain consistent, with the expert-guided method still performing best.
@@ 8 Conclusion
In this paper, we propose an expert-annotated knowledge base, evaluate its quality in the Similar Charge Distinction task, and apply it to the Legal Case Retrieval task. Our results demonstrate that expert annotations significantly enhance LLMs’ understanding of the Four-Element Theory. The four-element annotations, enriched with professional legal interpretations, provide strong support for LLMs’ reasoning capabilities. This approach can be extended to other legal AI tasks, such as legal document analysis and contract interpretation.
@@ 10 Limitations
As a limitation, this knowledge base focuses on the Four-Element Theory within the context of 155 crimes under Chinese Criminal Law. However, the four-level hierarchical pyramid annotation structure based on the legal interpretation system proposed in this work provides valuable insights for future expansion to other legal domains, as it represents a theoretical framework in the field of jurisprudence. The interpretative methods within the legal interpretation system, including textual, systematic, sociological, and doctrinal interpretations, are universally recognized in international law field and can be applied to different laws, countries, and legal systems.
@@ Acknowledgments
Appendix A Detailed Data Distribution for each Element
Figure 3: The average length distribution of total four elements annotated by experts.
@@ Appendix A Detailed Data Distribution for each Element
Figure 5: The average length distribution of total four elements generated by LLM.
@@ 2. Systematic Interpretation
An approach interpreting legal provisions through their position within the codified legal hierarchy and logical connections with related norms, maintaining the integrity of the legal system (aligned with Dworkin’s "law as integrity" theory).
@@ 4. Historical Interpretation
Interpretation based on legislative history materials including drafts, debates and official commentaries, while distinguishing original meaning from framers’ subjective intentions (as per Brest’s original understanding theory).
@@ 6. Sociological Interpretation
Interpretation evaluating social efficacy through empirical analysis of implementation effects, guided by Pound’s sociological jurisprudence principle that "law must be measured by its achieved results".
Appendix C Human Evaluation Guidance
The annotators included three postgraduate students specializing in criminal law and one master’s student in legal science and technology. The annotators scored independently, without knowledge of each other’s results. Before scoring, they were asked to read the descriptions and scoring guidelines (as shown in Table 5) for each evaluation dimension. In order to ensure the fairness of the evaluation, they do not know the source of each four elements, and even do not know that these four elements include those generated by LLMs.
@@ Appendix C Human Evaluation Guidance
Definition Whether there are errors in key elements Whether the four elements are complete Whether key elements and scenarios are emphasized Whether language and format are clear and standardized
@@
Appendix D Details for Similar Charge Disambiguation
@@
AP&DD Abuse of Power & Dereliction of Duty 1950 / 1938
Table 6: Distribution of charges in the GCI dataset. Cases denotes the number of cases in each category. Following Liu et al. (2021), for a case with both confusable charges, the prediction of any one of the charges is considered correct.
@@
The candidate charges and their corresponding four elements are as follows:
[Four Elements of Candidate Charges].
@@
Table 7: Prompt template for adding the Four-Element Theory and specific four elements of crime in charge disambiguation.
@@
To explore the effectiveness of notes-guided four elements in LLMs, we further consider other methods that introduced the Four-element theory into LLMs.
@@
MALR is a up to date multi-agent framework designed to enhance complex legal reasoning Yuan et al. (2024), enabling LLMs to autonomously decompose legal tasks and extract insights from legal rules. As its full implementation is not publicly available, we use the released code for the auto-planner module and implement the legal insight extraction following the specified steps and prompts, with necessary refinements. Experiments on the paper’s reported examples show that our implementation produces task decompositions and outputs largely consistent with the original results.
As shown in Table D, different methods differ in their prompts for generating and explaining the Four-Element Theory, but generally follow a similar process. For the SCD output, except for COT and MALR, which require reasoning processes and prediction results, all other methods only require the output of prediction results.
@@
Pre-task None None None LLM-generated four elements Expert-annotated four elements
@@
Candidate charges are as follows: #Candidate Charges The candidate charges and relevant legal articles are as follows: #Candidate Charges + #Articles Please analyze using the Four Elements Theory step by step: #details about each step. The candidate charges are as follows: #Candidate Charges The candidate charges and their corresponding four elements are as follows: #Four Elements of candidate charges. The four elements represent the four core factors of a charge. Compare the case facts to determine which charge’s four elements they align with, thereby identifying the charge.
Output format: #Format. Note: Only output the charge, no additional information.
Case facts: #Case Facts.
@@
Appendix E Baselines in Legal Case Retrieval
@@
BERTDevlin (2018) is a language model widely used in retrieval tasks. In this paper, we chose BERT-base-Chinese [^3] https://huggingface.co/google-bert/bert-base-chinese. Legal-BERT [^4] https://github.com/thunlp/OpenCLaPChalkidis et al. (2020) is a variant of BERT that is specifically trained on legal corpora. LawformerXiao et al. (2021)is a Chinese legal pre-trained model based on LongformerBeltagy et al. (2020), which is able to process long texts in the legal domain. ChatLaw-Text2Vec [^5] https://modelscope.cn/models/fengshan/ChatLaw-Text2VecCui et al. (2023) is a Chinese legal LLM trained on 936,727 legal cases for similarity calculation of legal-related texts. SAILERLi et al. (2023) is a structure-aware legal case retrieval model utilizing the structural information in legal case documents. GEARQin et al. (2024) is a generative retrieval framework that explicitly integrates judgment prediction with legal document retrieval in a sequence-to-sequence manner. Since the output of GEAR cannot directly evaluate NDCG, the official results under the same setting are directly referenced in this paper. LLM and Expert represent the results of retrieval using only the four elements.
Appendix F SCR results on the full LeCaRDv2 Dataset
@@
As presented in Table 9, we selected several representative methods based on sparse retrieval and dense retrieval for experiments on the full LeCaRDv2 dataset. All language models were not fine-tuned. The notes-guided FET method achieved the best performance among all language models, attaining top results in both R@500 and R@1000. The results indicate that the conclusions drawn from the full dataset are consistent with those from the subset, and the notes-guided method demonstrates strong performance.
@@
Appendix G A Case Study of LCR
Table 10 presents a case study on the Crime of Embezzlement. By comparing the four elements annotated by experts for the crime in JUREX-4E, the case-specific four elements generated directly by the LLM, and those generated by the LLM with expert four elements of charge as guidance, we can observe that:

### v2

@@ 8.3 Results
The LCR results are shown in Table 4, where we can observe that: (1) FET Enhances Retrieval. Integrating the FET improves retrieval performance across all metrics. For instance, BGE+FET(GPT-4o) improves MRR by 11.11%, and BGE+FET (Expert, GPT-4o) achieves an even larger gain of 11.89%, indicating that structured legal theory benefits retrieval quality. (2) Expert Knowledge is Important. Expert-guided case-FET consistently outperforms LLM-generated variants across both Qwen2.5-72b and GPT-4o backbones. For example, BGE+FET (Expert, GPT-4o) achieves higher Recall@30 (0.7967 vs. 0.7828) and MRR (0.2155 vs. 0.2140). The gap is even larger in the FET only setting (e.g., MRR 0.1624 vs. 0.1453 for GPT-4o), demonstrating that expert knowledge captures critical legal reasoning that LLMs may overlook.
We provide a case study in Appendix I. It illustrates that the expert-annotated FETs of charges provide practical judgment points and key narratives (e.g., the special subject of Embezzlement) that help the LLM focus on essential facts to analyze the case-FET.
9 Discussion
Model F&E Acc F&E F1 E&MPF Acc E&MPF F1 AP&DD Acc AP&DD F1 Avg. Acc Avg. F1
GPT-4o+FET4o 95.73 96.56 91.87 92.01 89.61 89.69 92.40 92.75
GPT-4o+FETCollaboration 95.99 96.56 91.51 91.61 91.02 90.95 92.84 93.04
GPT-4o+FETExpert 96.06 96.69 92.57 93.05 90.53 90.62 93.05 93.45
Table 5: Results of the collaboration study on Similar Charge Disambiguation (SCD).
To examine whether LLMs can effectively support expert annotation, we designed a preliminary collaboration pipeline that is roughly aligned with our annotation stages.
For each charge, the LLM first retrieved relevant statutory articles (the retrieval of legal sources in stages 1–2). It then extracted the top 10 factual keywords from cases cited by experts (the use of factual cues as in stage 3). Experts validated these keywords for each element, after which the LLM generated refined FET annotations by reasoning over the combined legal texts and factual cues. This pipeline yielded measurable improvements over unguided prompting (Table 5), illustrating the feasibility of expert-in-the-loop workflows.
Nevertheless, the study also highlights limitations. High-frequency keywords often lacked discriminative power for similar charges—for example, descriptors such as “beating” and “stabbing” occurred in both intentional injury and intentional homicide, reducing their utility. The refined FETs still fell short of the depth and precision of expert annotations. Future improvements could include decomposing FET generation into subtasks aligned with annotation stages and enriching each stage with more authoritative sources, thereby strengthening both coverage and normative awareness.
@@ 10 Conclusion
This paper presents JUREX-4E, an expert-annotated FET knowledge base built through a structured legal interpretation process and validated on downstream tasks. Grounded in widely accepted interpretive methods, our framework is adaptable across different branches of law and legal traditions, making it applicable beyond Chinese criminal law. Moreover, the structured approach to integrating expert domain knowledge may inspire applications in other areas where expert judgment is critical.
@@ 12 Limitations
Our current knowledge base is limited to 155 charges under Chinese Criminal Law due to the high cost of expert annotation. Future work will explore extending it to other legal domains and jurisdictions.
Another limitation lies in our current integration of factual and legal information. In the LCR task, although case facts are used to generate FETs, the FET only variant excludes the original case facts during retrieval, resulting in performance loss (e.g., MRR 0.1624 vs. 0.2155). This suggests that our current method remains coarse-grained, and more fine-grained fusion strategies, such as multi-agent coordination or retrieval-time integration, deserve future exploration.
@@ Acknowledgments
This work is supported in part by Beijing Science and Technology Program (Z231100007423011)
Appendix A Charge Selection and Detailed Data Distribution
Charge Selection: To systematically determine charge frequency, we analyzed the CAIL2018 dataset Xiao et al. (2018), which contains 2,676,075 criminal cases annotated with 183 criminal law articles and 202 criminal charges and a total of 3,010,000 criminal charges. Apart from few charges that have been merged or changed name, our dataset largely covers all criminal charges from CAIL2018 that have a frequency of over 3,000 (>0.099%) occurrences.
Length distribution for each element: Table 6.
Figure 3: The average length distribution of the total four elements annotated by experts.
@@ Appendix A Charge Selection and Detailed Data Distribution
Figure 5: The average length distribution of the total four elements generated by LLM.
@@ 2. Systematic Interpretation
An approach interpreting legal articles through their position within the codified legal hierarchy and logical connections with related norms, maintaining the integrity of the legal system (aligned with Dworkin’s "law as integrity" theory).
@@ 4. Historical Interpretation
Interpretation based on legislative history materials, including drafts, debates, and official commentaries, while distinguishing original meaning from framers’ subjective intentions (as per Brest’s original understanding theory).
@@ 6. Sociological Interpretation
Interpretation evaluating social efficacy through empirical analysis of implementation effects, guided by Pound’s sociological jurisprudence principle that “law must be measured by its achieved results”.
Appendix C Prompt for LLM-generated FET
See Table 6.
You are an expert in criminal law. Based on the given charge, please analyze it according to China’s criminal law and output the four elements of the charge in order, including:
- Object: The concretization of a certain abstract social interest. For example, the object of charges that infringe on personal rights is the right to life, while the object of property-related charges could be items such as mobile phones or wallets.
- Objective Aspect: The objective facts of the criminal act, including the key actions that trigger the charge (e.g., theft, robbery) and the consequences caused by the act (e.g., serious injury, death, property loss).
- Subject: Typically, the general subject of the charge, but in some cases, a specific subject is required (e.g., government officials in certain offenses).
- Subjective Aspect: The mental state of the perpetrator, such as intent or negligence.
Relevant Legal Articles: []
Please synthesize the above information to generate a refined set of four elements that represent the characteristics of the charge.
Output format: { "Crime": "", "Four-elements of the Crime": { "Crime Object": "", "Objective Aspect": "", "Subject": "", "Subjective Aspect": "" } }
Crime: []
Table 6: Prompt template for generating the four elements of a Crime using LLMs
Appendix D Details about Pilot Study
We selected candidate models from LawBench Fei et al. (2023) and LexEval Li et al. (2024c), which contain the broadest and most up-to-date evaluation of legal LLMs. From these, we chose top-performing models such as GPT-4, Qwen-14B-chat, and representative legal-specific LLMs.
For best performance, we used GPT-4o (the latest version of GPT-4 at that time) and Qwen-plus (a stronger commercial variant of Qwen2.5-72B. (Aliyun model-studio official site))
During implementation, we found that most legal LLMs were unavailable. The only stably accessible one was Farui (A leading legal LLM built on Qwen, (Aliyun model-studio official site, Tongyi Farui)), specifically the version “tongyifarui-890” from its official API.
To compare GPT-4o, Qwen-plus, and tongyifarui-890, we sampled 300 cases from our legal retrieval dataset and asked models to perform charge prediction, which is the pre-task for generating case-FETs.
(For each case in legal retrieval, the model was required to predict charges, so we can match charges’ expert-FETs, and use them to generate case-FET.)
This task involved all criminal charges, including multi-defendant and multi-charge scenarios, and requires models to predict charges from open text without a predefined list, making it a challenging legal task.
The result showed that GPT-4o (59.78%) > Qwen-plus (58.70%) >> tongyifarui-890 (21%). Given Farui’s poor performance, we did not include it in subsequent experiments.
We further evaluated GPT-4o and Qwen-plus based on their ability to generate case-FETs. The results showed that GPT-4o outperformed Qwen-plus (MRR 0.2140 vs. 0.2052). Considering both results, we adopted GPT-4o as our primary model in the paper.
Subsequently, in efforts to improve charge prediction for matching charges’ expert-FETs, we found that Qwen-plus performed better than GPT-4o when a charge list was provided (58.70%->80.43% vs. 59.78%->71.74%). Therefore, in this specific setting for charge prediction before retrieval, we used Qwen-plus.
For fair and reproducible presentation of results on specific downstream tasks (SCD and LCR), as mentioned in the main text, we present the results of open source Qwen2.5-72b.
Appendix E Human Evaluation Guidance
The annotators included three postgraduate students specializing in criminal law and one master’s student in legal science and technology. The annotators scored independently, without knowledge of each other’s results. Before scoring, they were asked to read the descriptions and scoring guidelines (as shown in Table 7) for each evaluation dimension. In order to ensure the fairness of the evaluation, they do not know the source of the four elements, and they do not know that the four elements include those generated by LLMs.
@@ Appendix E Human Evaluation Guidance
Definition Whether there are errors in key elements Whether the four-elements are complete Whether key elements and scenarios are emphasized Whether language and format are clear and standardized
@@
Appendix F Details for Similar Charge Disambiguation
@@
AP&$D$ Abuse of Power & Dereliction of Duty 1950 / 1938
Table 8: Distribution of charges in the GCI dataset. Cases denote the number of cases in each category. Following Liu et al. (2021), for a case with both confusable charges, the prediction of any one of the charges is considered correct.
@@
The candidate charges and their corresponding four-elements are as follows:
[four-elements of Candidate Charges].
@@
Table 9: Prompt template for adding the Four-Element Theory and specific four-elements of crime in charge disambiguation.
@@
To explore the effectiveness of expert-FETs, we further consider other methods that introduced the Four-element Theory into LLMs.
@@
MALR is an up-to-date multi-agent framework designed to enhance complex legal reasoning Yuan et al. (2024), enabling LLMs to autonomously decompose legal tasks and extract insights from legal rules. As its full implementation is not publicly available, we use the released code for the auto-planner module and implement the legal insight extraction following the specified steps and prompts, with necessary refinements. Experiments on the paper’s reported examples show that our implementation produces task decompositions and outputs largely consistent with the original results.
As shown in Table 10, different methods differ in their prompts for generating and explaining the Four-Element Theory, but generally follow a similar process. For the SCD output, except for COT and MALR, which require reasoning processes and prediction results, all other methods only require the output of prediction results.
@@
Pre-task None None None LLM-generated FETs Expert-annotated FETs
@@
Candidate charges are as follows: #Candidate Charges The candidate charges and relevant legal articles are as follows: #Candidate Charges + #Articles Please analyze using the four-elements Theory step by step: #details about each step. The candidate charges are as follows: #Candidate Charges The candidate charges and their corresponding four-elements are as follows: #four-elements of candidate charges. The four elements represent the four core factors of a charge. Compare the case facts to determine which charge’s four elements they align with, thereby identifying the charge.
Output format: #Format. Note: Only output the charge, no additional information. Case facts: #Case Facts.
@@
FET-LLM vs FET-Expert F&E E&MPF AP&$D$
FET-qwen vs FET-Expert 0.00215 0.00070 0.00509
FET-farui vs FET-Expert 0.00000 0.00126 0.00516
FET-4o+farui+qwen vs FET-Expert 0.00000 0.00996 0.03415
FET-4o vs FET-Expert 0.02246 0.00156 0.02251
Table 11: McNemar’s test results (p-values) comparing LLM-generated FET and expert-FET. Statistically significant improvements ($p<0.05$) are observed across all tasks.
Appendix G Baselines in Legal Case Retrieval
@@
BERT Devlin (2018) is a language model widely used in retrieval tasks. In this paper, we chose BERT-base-Chinese [^5] https://huggingface.co/google-bert/bert-base-chinese. Legal-BERT [^6] https://github.com/thunlp/OpenCLaP Chalkidis et al. (2020) is a variant of BERT that is specifically trained on legal corpora. Lawformer Xiao et al. (2021)is a Chinese legal pre-trained model based on Longformer Beltagy et al. (2020), which is able to process long texts in the legal domain. ChatLaw-Text2Vec [^7] https://modelscope.cn/models/fengshan/ChatLaw-Text2Vec Cui et al. (2023) is a Chinese legal LLM trained on 936,727 legal cases for similarity calculation of legal-related texts. SAILER Li et al. (2023) is a structure-aware legal case retrieval model utilizing the structural information in legal case documents.
Baseline results are provided in Table 12.
Model NDCG@10 NDCG@20 NDCG@30 R@1 R@5 R@10 R@20 R@30 MRR
BERT 0.1511 0.1794 0.1978 0.0199 0.0753 0.1299 0.2157 0.2579 0.1136
ChatLaw-Text2Vec 0.2049 0.2328 0.2745 0.0353 0.1306 0.1913 0.2684 0.3751 0.1285
BGE (case_fact only) 0.4737 0.5539 0.5937 0.0793 0.2945 0.4298 0.6500 0.7394 0.1926
BGE+FET (Qwen2.5) 0.5125 0.5858 0.6350 0.1104 0.2870 0.4653 0.6679 0.7836 0.2168
   FET only 0.3367 0.3971 0.4487 0.0622 0.2006 0.3279 0.4806 0.6037 0.1524
BGE+FET (Expert, Qwen2.5) 0.5295 0.5979 0.6416 0.1124 0.3122 0.4838 0.6791 0.7824 0.2206
   FET only 0.3354 0.4035 0.4541 0.0849 0.1923 0.3076 0.4839 0.6097 0.1606
BGE+FET (GPT-4o) 0.5139 0.5862 0.6291 0.0980 0.2967 0.4769 0.6802 0.7828 0.2140
   FET only 0.3583 0.4293 0.4798 0.0506 0.2240 0.3644 0.5383 0.6652 0.1453
BGE+FET (Expert, GPT-4o) 0.5211 0.5920 0.6379 0.1024 0.3049 0.4883 0.6885 0.7967 0.2155
   FET only 0.3766 0.4584 0.5111 0.0715 0.1894 0.3709 0.5891 0.7203 0.1624
Table 12: Performance on the Legal Charge Retrieval (LCR) task with baselines. Highest results are in bold. “FET only” indicates using the four-element descriptions without case facts.
To support reproducibility, we provide the full prompt templates used in our pipeline. Table 13 shows the prompt for charge prediction, and Table 14 presents the prompt used for generating four-element annotations in both BGE+FET(LLM) and BGE+FET(Expert, LLM).
Prompt 1: Charge Prediction
You are a legal expert specializing in criminal law. Based on the provided list of charges, determine which charges are applicable to the given case facts. Please note that you should only output the charge names, without any additional information. The charges must be selected from the provided list and should be separated by commas.
[Crime List]
[Case Facts]
Table 13: Prompt used for charge prediction.
Prompt 2: BGE+FET(LLM) and BGE+FET(Expert, LLM).
You are a legal expert specializing in criminal law. Based on Chinese criminal law knowledge, analyze the following case facts and provide the following information in sequence: 1. The four elements of the crime:
- Criminal Object: The tangible or intangible interests being infringed upon (e.g., personal rights such as life, or property rights such as money, vehicles).
- Objective Aspect: The objective facts of the criminal activity, including key actions (e.g., theft, robbery) and consequences (e.g., injury, death, loss).
- Criminal Subject: Typically general subjects; special subjects in certain crimes (e.g., government officials).
- Subjective Aspect: Whether the act was intentional or negligent.
2. Charge: Only output the specific crime name(s).
3. Relevant Legal Articles: Only output the article number(s) of the relevant laws.
Output format: JSON. For each crime involved in the case, provide a separate dictionary entry. [Output Sample]
"Crime 1": {
"Four elements": {
"Criminal Object": "Personal rights: the victim Wang’s right to life; Property rights: vehicle.",
"Objective Aspect": "The defendant Wu drove under the influence and collided with the victim Wang, causing Wang’s immediate death and vehicle damage.",
"Criminal Subject": "Defendant Wu, the driver.",
"Subjective Aspect": "Negligence"
"Charge": "Traffic Accident Crime",
"Relevant Legal Article": "Article 133"
"Crime 2": {
"Four elements": {
"Criminal Object": "Social management order: infringement on the state’s document management system; Property rights: forged documents and related items.",
"Objective Aspect": "Defendant 1 purchased equipment and materials to forge documents. Defendant 2 delivered the forged documents. Defendant 3 facilitated transactions via the internet, handling payments and document transfers.",
"Criminal Subject": "Multiple defendants, all individuals with full criminal responsibility.",
"Subjective Aspect": "Intentional"
"Charge": "Forgery, Alteration, or Sale of Official Documents, Certificates, and Seals of State Organs",
"Relevant Legal Article": "Article 280, Paragraph 1"
Table 14: Prompt for generating four-element annotations used in FETLLM and FETExpert_Guided.
Appendix H SCR results on the full LeCaRDv2 Dataset
@@
As presented in Table 15, we selected several representative methods based on sparse retrieval and dense retrieval for experiments on the full LeCaRDv2 dataset. All language models were not fine-tuned. The expert-guided FET method achieved the best performance among all language models, attaining top results in both R@500 and R@1000. The results indicate that the conclusions drawn from the full dataset are consistent with those from the subset, and the expert-guided method demonstrates strong performance.
@@
Appendix I A Case Study of LCR
Table 16 presents a case study on the Crime of Embezzlement. By comparing the expert-FETs for the charges in JUREX-4E, the case-FETs generated directly by the LLM, and those generated by the LLM with expert-FETs of charge as guidance, we can observe that: